OpenStudy (anonymous):

How to solve fractioned equations ? Like 1/2 = q+2/3

OpenStudy (anonymous):

Just like you would whole numbers. Your trying to find q? If so you would subtract 2/3 from both sides. 1/2- 2/3 \[\frac{1}{2 }-\frac{ 2 }{ 3 }=\]

OpenStudy (anonymous):

Find common den. So you answer would be \[-\frac{ 1 }{ 6 }\]